OpenStudy (anonymous):

PLEASE ANSWER How many tangent lines to the curve y = x/(x+1) pass through the point (1,2)? At which points do these tangent lines touch the curve?

OpenStudy (anonymous):

guess it didn't work want to try again?

OpenStudy (anonymous):

derivative is \[\frac{1}{(x+1)^2}\] and you know that the slope is \[\frac{y_2-y_1}{x_2-x_1}=\frac{\frac{x}{x+1}-2}{x-1}\] so now you have to solve \[\frac{1}{(x+1)^2}=\frac{\frac{x}{x+1}-2}{x-1}\]

OpenStudy (anonymous):

i am pretty sure this is right, even though the answer is rather ugly i get \[x=-2\pm\sqrt{3}\]

OpenStudy (anonymous):

show me how please